### v1

@@ Abstract
Recently, there has been interest in representing single graphs by multiple drawings; for example, using graph stories, storyplans, or uncrossed collections. In this paper, we apply this idea to orthogonal graph drawing. Due to the orthogonal drawing style, we focus on plane 4-graphs, that is, planar graphs of maximum degree 4 whose embedding is fixed. Our goal is to represent any plane 4-graph $G$ by an unbent collection, that is, a collection of orthogonal drawings of $G$ that adhere to the embedding of $G$ and ensure that each edge of $G$ is drawn without bends in at least one of the drawings. We investigate two objectives. First, we consider minimizing the number of drawings in an unbent collection. We prove that every plane 4-graph can be represented by a collection with at most three drawings, which is tight. We also give sufficient conditions for a graph to admit an unbent collection of size $2$. Second, we consider minimizing the total number of bends over all drawings in an unbent collection. We show that this problem is NP-hard and give a 3-approximation algorithm. For the special case of plane triconnected cubic graphs, we show how to compute minimum-bend collections in linear time.
@@
Graph drawing Orthogonal drawings Bend minimization
@@ 1 Introduction
A way to determine the quality of a drawing of a graph is its readability, which, informally speaking, is the extent to which a drawing conveys the structure of the graph it represents. More formally, readability translates into a set of goals that a graph drawing algorithm must optimize. For example, in the context of orthogonal graph drawing, the typical objectives are crossing minimization and bend minimization. Note that these objectives may be opposing each other since reducing the number of crossings may lead to adding bends along the edges, and vice versa. In fact, minimizing the number of bends in a planar (i.e., crossing-free) orthogonal drawing may require some edges to have linearly many bends [16].
@@
A recent research direction is to represent a graph by a collection or sequence of different drawings, each of which optimizes the readability of a particular subgraph, possibly at the expenses of the rest. One contribution in this direction is the so-called storyplan approach, where a graph is represented by a sequence of drawings, called frames, each of which only shows a restricted (e.g., planar) subgraph. To see the whole (complicated) graph means to watch all (relatively simple) frames [2, 3, 6]. Another approach is to visualize the graph by means of a collection of drawings, each of which includes all vertices and edges, but to optimize the drawings only with respect to specific subgraphs. Hliněný and Masařík [9] adopted this idea and introduced the problem of representing a non-planar graph $G$ via multiple drawings such that $G$ is fully represented in each drawing and each edge of $G$ is not crossed in at least one of the drawings. They called such a collection an uncrossed collection. The size of a collection is the number of drawings it contains. Hliněný and Masařík studied two optimization problems. For a given graph $G$, the first problem called the uncrossed number is to find a minimum-size uncrossed collection of $G$. The second problem is to find an uncrossed collection that minimizes the total number of crossings over all drawings. Recently, the uncrossed number was further studied in [1]. It has been exactly determined for complete and complete bipartite graphs. Additionally, the authors presented a strong lower bound on general graphs in terms of the number of vertices and edges of $G$.
@@
We revisit the classical problem of computing a planar orthogonal layout with the minimum number of bends from the “collection perspective” of Hliněný and Masařík. Namely, we study the following problem: Given a plane 4-graph $G$ (i.e., a planar graph with vertex degree at most $4$ and with a fixed planar embedding), compute a set of planar orthogonal drawings of $G$ such that every edge of $G$ is straight (i.e., has no bend) in at least one drawing of the collection. Analogously to the terminology of Hliněný and Masařík, we call such a collection an unbent collection. We study both the problem of finding a minimum-size unbent collection and the problem of finding an unbent collection that minimizes the total number of bends. Refer to Figure 1 for an example.
Figure 1: (a) A minimum-bend orthogonal drawing of $K_{4}$. (b) An unbent collection for $K_{4}$ with two drawings showing that $\operatorname{un}(K_{4})=2$ and $\operatorname{tbn}(K_{4})\leq 12$ (in fact, $\operatorname{tbn}(K_{4})=12$); see Definition 1. Red edges are straight, whereas blue edges are bent.
Specifically, we show the following. For every plane 4-graph, an unbent collection of size 3 always exists and can be found in almost linear time; see Theorem 2.1. We also give a necessary and sufficient condition for graphs to have an unbent collection of size 2. The condition is based on a new concept of balanced 2-edge-coloring that we introduce in Section 2. Then, we show that it is NP-hard to decide, given a plane 4-graph $G$ and a positive integer $k$, whether $G$ admits an unbent collection with at most $k$ bends; see Section 3. Note that if the embedding of $G$ is not fixed, it is even NP-hard to decide whether $G$ admits a (single) drawing without bends [8]. Next, we give a 3-approximation algorithm for minimizing the number of bends in an unbent collection. It relies on Tamassia’s flow network [15], combined with local modifications and the insight that every 4-graph has star arboricity at most 4.
@@
For statements marked with a clickable ()($\star$)( ⋆ ), we give a proof in the appendix.
@@
Definition 1
Let $G$ be a plane 4-graph. A set of orthogonal drawings of $G$ (all adhering to the embedding of $G$) is called an unbent collection of $G$ if every edge of $G$ is drawn without a bend (i.e., straight) in at least one drawing of the set. The unbent number of $G$, denoted $\operatorname{un}(G)$, is the size of a smallest unbent collection of $G$. The unbent bend number of $G$, denoted $\operatorname{tbn}(G)$, is the minimal total number of bends, taken over all unbent collections of $G$.
Given an $n$-vertex plane 4-graph $G$, Tamassia [15] defined a flow network $N_{G}$ that has a node for each vertex and for each face of $G$, an arc from each vertex to each incident face, and two arcs for each pair of adjacent faces of $G$, one in each direction across the corresponding edge of $G$. A unit of flow on an edge from node $a$ to node $b$ represents a 90 degrees angle. Flow originates in vertices and faces of degree 3 and terminates in faces of degree greater than 5 and the external face. For the sake of presentation, we invert the flow direction in this paper. More precisely, each vertex has a demand of $4$, each inner face $f$ has a demand of $4-\deg(f)$, and the external face $f_{\operatorname{ext}}$ has a demand of $-4-\deg(f_{\operatorname{ext}})$. Flow in $N_{G}$ has non-zero cost only for arcs between vertices and faces; otherwise, the cost is one for each unit of flow. The total cost equals the number of bends on the edges of $G$ in the resulting orthogonal representation [15], which refines the given embedding by specifying angles and bends (but not the lengths of edges). Such a representation can then be turned into an orthogonal drawing in time linear in the number of vertices and bends [15].
Note that there is an algorithm that computes a minimum-cost flow in $O(m^{1+o(1)}\log U\log C)$ time [4], where $m$ is the number of edges, $U$ is the maximum edge capacity, and $C$ is the maximum absolute edge cost. In $N_{G}$, $m\in O(n)$, $U\in O(n)$, and $C\in O(1)$. Since $\log n=n^{\log\log n/\log n}\in n^{o(1)}$, the algorithm runs in $O(n^{1+o(1)})$ time on input $N_{G}$.
@@ 2 Minimizing the Size of an Unbent Collection
Proposition 1
If the edges of a plane 4-graph $G$ can be covered by $k$ edge-disjoint forests, then $\operatorname{un}(G)\leq k$. Given $G$ and such a set of forests, an unbent collection realizing $\operatorname{un}(G)$ can be computed in $O(kn^{1+o(1)})$ time, where $n$ is the number of vertices of $G$.
Proof
For $i\in[k]$, let $F_{i}$ be the $i$-th given forest. We construct Tamassia’s flow network for $G$. We then make $k$ copies of the network. In the $i$-th copy, we delete the arcs corresponding to the edges of $F_{i}$. Solving these $k$ flow networks individually produces $k$ orthogonal drawings. Each edge $e$ of $G$ is drawn straight in at least one of these drawings, namely in the $i$-th drawing if $e$ lies in $F_{i}$. Note that there always exists a solution because, as we remove only edges forming forests but no cycles, the nodes in the flow network representing the faces still form a strongly connected component. For each $i\in[k]$, constructing the $i$-th flow network and removing the links corresponding to edges in $F_{i}$ can be done in linear time, and each flow network can be solved in $O(n^{1+o(1)})$ time [4].
Proposition 1 and Schnyder woods [14] yield an upper bound on $\operatorname{un}(G)$.
Theorem 2.1
@@
Proof
Extend $G$ to a plane triangulation $G^{+}$, and decompose $E(G^{+})$ into three disjoint trees in linear time by Schnyder’s algorithm [14]. Ignoring the edges from $G^{+}\setminus G$ leaves three edge-disjoint forests, to which we apply Proposition 1.
Theorem 2.1 tells us that for any planar graph $G$ of maximum degree $4$, we have $\operatorname{un}(G)\leq 3$. We will later show that this bound is in fact tight, by constructing an infinite family of graphs with $\operatorname{un}(G)=3$. Using Tamassia’s flow network [15], we can decide efficiently whether $\operatorname{un}(G)=1$. Thus, the real challenge is to decide whether two drawings are sufficient or three are necessary. We give a partial answer to this question below.
Proposition 2
Let $G$ be a plane 4-graph. If for every subset $V^{\prime}\subseteq V(G)$, the induced subgraph $G[V^{\prime}]$ hast at most $22|V^{\prime}|-$ 2 edges, then $\operatorname{un}(G)\leq 2$.
Proof
According to Nash-Williams’ theorem [11], a graph $H$ can be decomposed into $k$ edge-disjoint forests if and only if, for every $U^{\prime}\subseteq V(H)$, the induced graph $H[U^{\prime}]$ has at most $k(|U^{\prime}|-1)$ edges. Applying this and Proposition 1 to $G$ and $k=2$ yields $\operatorname{un}(G)\leq 2$.
Testing the sufficient criterion presented in Proposition 2 naively takes exponential time. So, we next present a way to decompose $G$ into relevant subgraphs.
@@
Proof (sketch)
Essentially, we decompose $G$ into induced subgraphs such that each such subgraph $G^{\prime}$ has at least $12|V(G^{\prime})|-$ 1 edges. (The other parts of $G$ can be ignored because, due to Proposition 2, they admit an unbent collection of size 2 anyway.) Since $G$ has maximum degree 4 and $G^{\prime}$ misses at most one edge to be 4-regular, $G^{\prime}$ is connected to $G\setminus G^{\prime}$ with at most two edges. Hence, we find all 1- and 2-edge cuts in linear time, which decompose the graph into smaller graphs. If and only if for each such small graph $G^{\prime}$, $\operatorname{un}(G^{\prime})\leq 2$ holds, then $\operatorname{un}(G)\leq 2$ also holds. In the smaller graph, we keep the edges of the cut, which would end at leaves (i.e., vertices of degree 1). To still match $|E(G^{\prime})|\geq 2|V(G^{\prime})|-1$ for each $G^{\prime}$, we attach the gadget $H$ shown in Fig. 2 to each such leaf.
Figure 2: (a) The gadget $H$ with $12n-$ 1 edges used to replace a degree-1 vertex in the proof of Lemma 1. (b) An unbent collection of size 2 for $H$.
Lemma 1 tells us that we should focus on graphs with at least $12n-$ 1 edges (a 4-graph has at most $2n$ edges, so these are the “densest” plane 4-graphs) as the calculation of the unbent number of all graphs with lower density boils down to these graphs. It is natural to ask whether graphs with $\operatorname{un}(G)=3$ even exist, which we answer positively later. If $G$ has at least $12n-$ 1 edges, then Proposition 1 cannot be applied to obtain $\operatorname{un}(G)\leq 2$ as every 2-edge-coloring will contain a cycle in at least one color. So, we next characterize graphs of unbent number at most 2 via 2-edge-colorings with specific properties.
@@
With this characterization, we can show that graphs with unbent number 3 exist. We define the flower graph with $k$ petals, $F_{k}$, to be the graph obtained from two cycles of length $k$, $C=(v_{1},\dots,v_{k})$ and $C^{\prime}=(v^{\prime}_{1},\dots v^{\prime}_{k})$ by adding edges $\{v_{i},v^{\prime}_{i}\}$ and $\{v_{i},v^{\prime}_{i+1}\}$ for each $i\in[k]$. For an illustration, see Fig. 4. Note that $F_{3}$ is the octahedron.
@@
Proposition 4 shows that $K_{5}$ minus an edge and the octahedron minus an edge have unbent number 3 if the large face is the external face; see Fig. 3(d). Thus, plane 4-graphs that contain one of those two as induced (plane) subgraph also has unbent number 3; see Fig. 3(e). It also shows $\operatorname{un}(F_{k})=3$ for $k\in[3,7]$, and that the 4-regular graph in Fig. 3(c) has unbent number 3. We remark that $\operatorname{un}(F_{k})=2$ for $k\geq 8$; see Fig. 3(b).
@@
Proof (sketch)
Suppose $\operatorname{un}(G)\leq 2$. Then, by Proposition 3, $G$ admits a balanced 2-edge-coloring. Let us first assume condition (i). The edge of one color must induce a cycle since $m=2n-1$. This cycle, however, encloses only triangles and thus cannot be balanced. Hence, $\operatorname{un}(G)=3$.
Now assume condition (ii). First suppose that one color induces two cycles $c_{1}$ and $c_{2}$ and let $F_{1}$ and $F_{2}$ be the sets of faces enclosed by $c_{1}$ and $c_{2}$, respectively. Because both cycles are balanced, both $F_{1}$ and $F_{2}$ contain $f$. The faces in the symmetric difference $F_{1}\triangle F_{2}$ are only triangles that are enclosed by at least one monochromatic cycle, which cannot be balanced. Hence, the edges of each color induce a spanning graph with exactly one (balanced) cycle containing $f$. This is not possible if $f$ shares an edge with the external face. Otherwise, let $k$ be the number of edges of $f$. Then each edge $e$ of $f$ is incident to another triangle that is contained in the the cycle of the color not assigned to $e$. This results in at least $k$ triangles inside of both monochromatic cycles, while, for $k\leq 7$, there should be only $2(\deg(f)-4)=2k-8<k$ triangular faces to account for the demand of $f$.

### v2

@@ Abstract
Recently, there has been interest in representing single graphs by multiple drawings; for example, using graph stories, storyplans, or uncrossed collections. In this paper, we apply this idea to orthogonal graph drawing. Due to the orthogonal drawing style, we focus on 4-graphs, that is, graphs of maximum degree 4. We restrict ourselves to plane graphs, that is, planar graphs whose embedding is fixed. Our goal is to represent any plane 4-graph $G$ by an unbent collection, that is, a collection of orthogonal drawings of $G$ that adhere to the embedding of $G$ and ensure that each edge of $G$ is drawn without bends in at least one of the drawings. We investigate two objectives. First, we consider minimizing the number of drawings in an unbent collection. We prove that every plane 4-graph can be represented by a collection with at most three drawings, which is tight. We also give necessary and sufficient conditions for a graph to admit an unbent collection of size $2$. Second, we consider minimizing the total number of bends over all drawings in an unbent collection. We show that this problem is NP-hard and give a 3-approximation algorithm. For the special case of plane triconnected cubic graphs, we show how to compute minimum-bend collections in linear time.
@@
Orthogonal drawings, Bend minimization, Unbent collection
@@ 1 Introduction
A way to determine the quality of a drawing of a graph is its readability, which, informally speaking, is the extent to which a drawing conveys the structure of the graph it represents. More formally, readability translates into a set of goals that a graph drawing algorithm must optimize. For example, in the context of orthogonal graph drawing, the typical objectives are crossing minimization and bend minimization. These objectives typically oppose each other: reducing the number of crossings may lead to adding bends along the edges, and vice versa.
@@
A recent research direction is to represent a graph by a collection or sequence of different drawings, each of which optimizes the readability of a particular subgraph, possibly at the expense of the rest. One contribution in this direction is the so-called storyplan approach, where a graph is represented by a sequence of drawings, called frames, each of which only shows a restricted (e.g., planar) subgraph. To see the whole (complicated) graph means to watch all (relatively simple) frames [4, 7]. Another contribution are graph stories, where each frame contains the same number of vertices, vertices are shown in an interval of consecutive frames, and, for each vertex, the index of the first frame when it appears is given as input. As in a storyplan, it is required that the drawing of an edge and the position of a vertex is the same in every frame where the edge or the vertex appears. The task is to compute graph stories with specific properties (e.g., crossing-free frames [3]). Another approach is to visualize the graph by means of a collection of drawings, each of which includes all vertices and edges, but to optimize the drawings only with respect to specific subgraphs. Hliněný and Masařík [9] adopted this idea and introduced the problem of representing a non-planar graph $G$ via multiple drawings such that $G$ is fully represented in each drawing and each edge of $G$ is not crossed in at least one of the drawings. They called such a collection an uncrossed collection. The size of a collection is the number of drawings it contains. Hliněný and Masařík studied two optimization problems. For a given graph $G$, the first problem called the uncrossed number is to find a minimum-size uncrossed collection of $G$. The second problem is to find an uncrossed collection that minimizes the total number of crossings over all drawings, for which they showed an NP-hardness and an FPT algorithm parameterized by the solution size. Recently, the authors of [2] determined the uncrossed number of complete and complete bipartite graphs exactly. They also gave a lower bound for the uncrossed number of general graphs in terms of vertex and edge numbers.
@@
We revisit the classical problem of computing a planar orthogonal layout with the minimum number of bends from the “collection perspective” of Hliněný and Masařík. Namely, we study the following problem: Given a plane 4-graph $G$ (i.e., a planar graph with vertex degree at most $4$ and with a fixed planar embedding), compute a set of planar orthogonal drawings of $G$ such that every edge of $G$ is straight (i.e., has no bend) in at least one drawing of the collection. Analogously to the terminology of Hliněný and Masařík, we call such a collection an unbent collection. We study two problems. For a given graph $G$, we want to find an unbent collection of minimum size, $\operatorname{un}(G)$, and an unbent collection that minimizes the total number of bends, $\operatorname{tbn}(G)$. Refer to Figure 1.
Figure 1: (a) A minimum-bend orthogonal drawing of $K_{4}$. (b) An unbent collection for $K_{4}$ with two drawings showing that $\operatorname{un}(K_{4})=2$ and $\operatorname{tbn}(K_{4})\leq 12$ (in fact, $\operatorname{tbn}(K_{4})=12$); see Definition 1.1. Red edges are straight, whereas blue edges are bent.
Specifically, we show the following. For every plane 4-graph, an unbent collection of size 3 always exists (this is tight) and can be found in almost-linear time; see Theorem 2.3. We also give a necessary and sufficient condition for graphs to have an unbent collection of size 2. The condition is based on a new concept of balanced 2-edge-coloring that we introduce in Section 2.
Then, we show that it is NP-hard to decide, given a plane 4-graph $G$ and a positive integer $k$, whether $G$ admits an unbent collection with at most $k$ bends; see Section 3. Note that if the embedding of $G$ is not fixed, it is even NP-hard to decide whether $G$ admits a (single) drawing without bends [8]. Next, we give a 3-approximation algorithm for minimizing the number of bends in an unbent collection. It relies on Tamassia’s flow network [15], combined with local modifications and the insight that every 4-graph has star arboricity at most 4.
@@
We prove statements marked with a clickable ()($\star$)( ⋆ ) in the appendix.
@@
Definition 1.1.
Let $G$ be a plane 4-graph. A set of orthogonal drawings of $G$ (all adhering to the embedding of $G$) is called an unbent collection of $G$ if every edge of $G$ is drawn without a bend (i.e., straight) in at least one drawing of the set. The unbent number of $G$, denoted $\operatorname{un}(G)$, is the size of a smallest unbent collection of $G$. The unbent bend number of $G$, denoted $\operatorname{tbn}(G)$, is the minimum total number of bends, taken over all unbent collections of $G$.
Given a plane 4-graph $G$, Tamassia [15] defined a flow network $N_{G}$ that has a node for each vertex and for each face of $G$, an arc from every face to each incident vertex, and two parallel arcs (one in each direction) for each edge $e$ of $G$ connecting the faces incident to $e$. A unit of flow in $N_{G}$ represents a $90^{\circ}$ angle at a vertex of $G$ or a bend on an edge of $G$. For the sake of presentation, we invert the usual flow direction in this paper. Then, each vertex has a demand of $4$, each inner face $f$ has a demand of $4-\deg(f)$, and the external face $f_{\operatorname{ext}}$ has a demand of $-4-\deg(f_{\operatorname{ext}})$. In the case of negative demand (i.e., flow excess), we also speak of supply (of flow). Faces of degree greater than 4 and the external face supply flow, while vertices and faces of degree 3 demand flow. Flow in $N_{G}$ has non-zero cost only for arcs between faces and vertices; otherwise, the cost is one for each unit of flow. The total cost equals the number of bends on the edges of $G$ in the resulting orthogonal representation [15], which refines the given embedding by specifying angles and bends (but not the lengths of edges). Such a representation can then be turned into an orthogonal drawing in time linear in the number of vertices and bends [15].
Recently, a minimum-cost flow algorithm that runs in $O(m^{1+o(1)}\log U\log C)$ time was presented [16], where $m$ is the number of edges, $U$ is the maximum edge capacity, and $C$ is the maximum absolute edge cost in the flow network. In $N_{G}$, $C\in O(1)$, $m\in O(n)$, and $U\in O(n)$, where $n=|V(G)|$. Since $\log n=n^{\log\log n/\log n}\in n^{o(1)}$, the algorithm runs in $O(n^{1+o(1)})$ time on input $N_{G}$.
@@ 2 Minimizing the Size of an Unbent Collection
Proposition 2.1.
If a plane 4-graph $G$ can be partitioned into $k$ forests, then $\operatorname{un}(G)\leq k$. Given $G$ and such a set of forests, an unbent collection of size $k$ can be computed in $O(kn^{1+o(1)})$ time, where $n$ is the number of vertices of $G$.
Proof 2.2.
For $i\in[k]$, let $F_{i}$ be the $i$-th given forest. We construct Tamassia’s flow network for $G$. We then make $k$ copies of the network. In the $i$-th copy, for each edge $e\in F_{i}$, we delete the two arcs that connect the two nodes representing the incident faces of $e$. Since $F_{i}$ contains no cycle, note that the nodes in the flow network representing the faces still form a strongly connected component, hence there always exists a solution. Solving these $k$ flow networks individually produces $k$ orthogonal drawings. Each edge $e$ of $G$ is drawn straight in at least one of these drawings, namely in the $i$-th drawing if $e$ lies in $F_{i}$. For each ${i\in[k]}$, constructing the $i$-th flow network and removing the links corresponding to edges in $F_{i}$ can be done in linear time, and each flow network can be solved in $O(n^{1+o(1)})$ time [16].
Proposition 2.1 and Schnyder woods [14] yield an upper bound on $\operatorname{un}(G)$.
Theorem 2.3.
@@
Proof 2.4.
Extend $G$ to a plane triangulation $G^{+}$, and decompose $E(G^{+})$ into three disjoint trees in linear time by Schnyder’s algorithm [14]. Ignoring the edges from $G^{+}\setminus G$ leaves three edge-disjoint forests, to which we apply Proposition 2.1.
Theorem 2.3 tells us that for any planar graph $G$ of maximum degree $4$, we have $\operatorname{un}(G)\leq 3$. We will later show that this bound is in fact tight, by constructing an infinite family of graphs with unbent number $3$. Using Tamassia’s flow network [15], we can decide efficiently whether $\operatorname{un}(G)=1$. Thus, the real challenge is to decide whether two drawings are sufficient or three are necessary. We give a partial answer to this question below.
Proposition 2.5.
Let $G$ be a plane 4-graph. If, for every subset $V^{\prime}\subseteq V(G)$, the induced subgraph $G[V^{\prime}]$ has at most $22|V^{\prime}|-$ 2 edges, then $\operatorname{un}(G)\leq 2$.
Proof 2.6.
According to Nash-Williams’ theorem [11], a graph $H$ can be partitioned into $k$ forests if and only if, for every $U^{\prime}\subseteq V(H)$, the induced graph $H[U^{\prime}]$ has at most $k\cdot(|U^{\prime}|-1)$ edges. Applying Nash-Williams’ theorem and Proposition 2.1 to $G$ and $k=2$ yields $\operatorname{un}(G)\leq 2$.
Testing the sufficient criterion presented in Proposition 2.5 naively takes exponential time. So, we next present a way to decompose $G$ into relevant subgraphs.
@@
Proof 2.8 (Proof sketch).
Essentially, we decompose $G$ into induced subgraphs such that each such subgraph $G^{\prime}$ has at least $12|V(G^{\prime})|-$ 1 edges. (The other parts of $G$ can be ignored because, due to Proposition 2.5, they admit an unbent collection of size 2 anyway.) Since $G$ has maximum degree 4 and $G^{\prime}$ misses at most one edge to be 4-regular, $G^{\prime}$ is connected to $G\setminus G^{\prime}$ with at most two edges. Hence, we find all 1- and 2-edge cuts in linear time, which decompose the graph into smaller graphs. In each smaller graph, we keep the edges of the cuts, which would end at leaves (i.e., vertices of degree 1). To still match $|E(G^{\prime})|\geq 2|V(G^{\prime})|-1$ for each smaller graph $G^{\prime}$, we attach the gadget $H$ shown in Figure 2 to each such leaf. If $\operatorname{un}(G^{\prime})\leq 2$ holds for each $G^{\prime}$, we obtain an unbent collection of size $2$ of $G$ by merging the smaller graphs at the leaves of the corresponding cuts. Since each cut has size at most $2$, we can keep one edge straight in each of the two drawings and add as many bends as necessary to the other edge.
Figure 2: (a) The gadget $H$ with $12n-$ 1 edges used to replace a degree-1 vertex in the proof of Lemma 2.7. (b) An unbent collection of size 2 for $H$.
Lemma 2.7 tells us that we should focus on graphs with at least $12n-$ 1 edges (a 4-graph has at most $2n$ edges, so these are the “densest” plane 4-graphs) as the calculation of the unbent number of all graphs with lower density boils down to these graphs. It is natural to ask whether graphs with $\operatorname{un}(G)=3$ even exist, which we answer positively later. If $G$ has at least $12n-$ 1 edges, then Proposition 2.1 cannot be applied to obtain $\operatorname{un}(G)\leq 2$ as every 2-edge-coloring will contain a cycle in at least one color. So, we next characterize graphs of unbent number at most 2 via 2-edge-colorings with specific properties.
@@
With this characterization, we can show that graphs with unbent number 3 exist. We define the flower graph with $k$ petals, $F_{k}$, to be the graph obtained from two cycles of length $k$, $C=(v_{1},\dots,v_{k})$ and $C^{\prime}=(v^{\prime}_{1},\dots,v^{\prime}_{k})$ by adding edges $\{v_{i},v^{\prime}_{i}\}$ and $\{v_{i},v^{\prime}_{i+1}\}$ for each $i\in[k]$. For an illustration, see Figure 4. Note that $F_{3}$ is the octahedron.
@@
The following Proposition 2.10 shows that $K_{5}$ minus an edge and the octahedron minus an edge have unbent number 3 if the large face is the external face; see Figure 3(d). Thus, plane 4-graphs that contain one of those two as an induced (plane) subgraph also have unbent number 3; see Figure 3(e). It also shows ${\operatorname{un}(F_{k})=3}$ for $k\in[3,7]$, and that the 4-regular graph in Figure 3(c) has unbent number 3. We remark that $\operatorname{un}(F_{k})=2$ for $k\geq 8$; see Figure 3(b).
@@
Proof 2.11 (Proof sketch).
Suppose $\operatorname{un}(G)\leq 2$. Then, by Proposition 2.9, $G$ admits a balanced 2-edge-coloring. Let us first assume condition (i). The edges of one color must induce a cycle since $m=2n-1$. This cycle, however, encloses only triangles and thus cannot be balanced. Hence, $\operatorname{un}(G)=3$.
Now assume condition (ii). First suppose that one color induces two cycles $c_{1}$ and $c_{2}$ and let $F_{1}$ and $F_{2}$ be the sets of faces enclosed by $c_{1}$ and $c_{2}$, respectively. Because both cycles are balanced, both $F_{1}$ and $F_{2}$ contain $f$. The faces in the symmetric difference $F_{1}\triangle F_{2}$ are only triangles that are enclosed by at least one monochromatic cycle, which cannot be balanced. Hence, the edges of each color induce a spanning graph with exactly one (balanced) cycle containing $f$. This is not possible if $f$ shares an edge with the external face. Otherwise, let $k$ be the number of edges of $f$. Then each edge $e$ of $f$ is incident to another triangle that is contained in the cycle of the color not assigned to $e$. This results in at least $k$ triangles inside of both monochromatic cycles, while, for $k\leq 7$, there should be only $2(\deg(f)-4)=2k-8<k$ triangular faces to account for the demand of $f$.